# Multi-scale Topology Optimization using Neural Networks

### v1

@@ Abstract
A long-standing challenge is designing multi-scale structures with good connectivity between cells while optimizing each cell to reach close to the theoretical performance limit. We propose a new method for direct multi-scale topology optimization using neural networks. Our approach focuses on inverse homogenization that seamlessly maintains compatibility across neighboring microstructure cells. Our approach consists of a topology neural network that optimizes the microstructure shape and distribution across the design domain as a continuous field. Each microstructure cell is optimized based on a specified elasticity tensor that also accommodates in-plane rotations. The neural network takes as input the local coordinates within a cell to represent the density distribution within a cell, as well as the global coordinates of each cell to design spatially varying microstructure cells. As such, our approach models an n-dimensional multi-scale optimization problem as a 2n-dimensional inverse homogenization problem using neural networks. During the inverse homogenization of each unit cell, we extend the boundary of each cell by scaling the input coordinates such that the boundaries of neighboring cells are combined. Inverse homogenization on the combined cell improves connectivity. We demonstrate our method through the design and optimization of graded multi-scale structures.
@@ 1 INTRODUCTION
There has been a recent increase in machine learning-driven topology and multi-scale structures optimization approaches, particularly using neural networks for performing structural optimization. Both data-driven [1, 2, 3, 4, 5] and direct training-based [6, 7, 8, 9, 10, 11] approaches have been explored. Data-driven approaches typically require large training sets and long training times. They perform instant optimal topology generation during inference time. Direct training approaches use the neural network to represent the density field of a single or a group of designs. Prior works on direct training approaches focus on using the neural network with spatial coordinates as input and density as the output to represent the geometry [9, 10, 6]. Inverse homogenization refers to the optimization of microstructure with the desired property [12]. Sridhara et al. [10] applied direct neural network-based inverse homogenization in designing single microstructure cells. Unfortunately, no direct neural network-based optimization for multi-scale structures has been developed. One key challenge from microstructure design to multi-scale structure is ensuring good connectivity between cells [12]. Connecting individual optimized microstructure cells into a multi-scale structure is called de-homogenization [12]. Cell rotation is included in de-homogenization in which cells can perform very close to the theoretical limit[12]. De-homogenization is a post-processing step to obtain a well-connected multi-scale design introduced by Pantz and Trabelsi [13]. During inverse homogenization, connectivity can also be encouraged where Garner et al. [14] combined adjacent cells and Liu et al. [15] formulated a constraint on the boundary of each cell. However, these previous methods did not incorporate rotation during inverse homogenization. On the other hand, de-homogenization relies on post-processing to add rotation.
@@
In this work, we aim to use a neural network to represent the design of a multi-scale structure such that connectivity and rotation and be enforced. We configure a neural network to take in both local and global spatial coordinates to represent multi-scale structural designs. We focus on 2D multi-scale structures, hence the coordinates are two-dimensional. The local coordinates represent the fine-scale microstructure. The global coordinates correspond with the coarse-scale cell location. Given the local and global coordinate pairs as input, the neural network outputs the density for that specific location as a value between 0 and 1. During training111Here “training” refers to optimization and back-propagation. Training has been used in previous related works [6, 11] but this is different from data-driven machine learning, the microstructure for each cell is optimized to maximize or minimize a specific material property, derived from the elasticity tensor. This is often referred to as inverse homogenization [16]. We extend the unit cell boundary to its neighboring cells during inverse homogenization to encourage connectivity between cells. Since the inputs to the neural network are spatial coordinates, transformation matrices, for example, scaling, and rotation, can be applied to the coordinates. We apply rotation matrices to the local input coordinates such that the cell is oriented similarly to de-homogenization[13]. The difference is that in our approach the orientation of the cell is co-optimized during the training of the neural network instead of a post-processing step.
@@
accommodate rotation of cells with neural network geometry representation
@@ 2 RELATED WORK
Multi-scale topology optimization: Multi-scale structures offer the potential to deliver outstanding performance, being inherently lightweight, durable, and versatile. Bendsøe and Kikuchi [17] introduced the homogenization approach for topology optimization. The SIMP method [18, 19] considers the relative material density in each element of the Finite Element (FE) mesh as design variables, allowing for a simpler interpretation and optimized designs with more clearly defined features. A comprehensive review by Wu et al. [12] discusses the existing approaches of multi-scale topology optimization and analysis of their strength and applications. Research that targets improving the compatibility of multi-scale structures includes Garner et al. [14] where adjacent cells were combined when inverse homogenization is performed such that compatibility can be improved. However, adding neighboring cells to the optimization will increase the computation time for FE analysis. We instead extend the cell boundary into neighboring cells where the FE resolution remains the same. Liu et al. [15] formulated a constraint on the boundary of each cell such that given n types of desired microstructure, all these n types share the same boundary. This allows the optimized cell designs to be mutually compatible. One possible limitation is that the total types of cell designs are limited and as the number of cell designs increases, the restriction on the boundary can be increased. In this work, we establish connectivity between neighboring cells, where the limitation on geometry is less. Connectivity is also important when designing multi-scale structures given a material exemplar. Senhora et al. [20] use spinodal architected materials to place material along principal stress directions. A library of microstructure can be pre-generated and connectivity between cells can be improved through mapping [21]. However, the microstructure performance gamut can be limited by the generated dataset.
@@ 3.1 Neural network-based topology representation
The topology network (Figure 1), learns a density field differently as compared to typical topology optimization which represents the density field as a finite element mesh. The topology neural network takes in local domain coordinates , as well as the global coordinate for each cell. The global coordinates get concatenated with the local coordinates to form the input to the topology network, . The local coordinates are normalized between to . The global coordinates are normalized between between to along the longest axis. Given the local and global coordinates, the topology network outputs the density value at each coordinate point. The domain coordinates represent the center of each element in the design domain. During topology optimization, a batch of domain coordinates that correspond to the mesh grid and the corresponding strain energy field is fed into the topology network. The output is then sent to the FE solver. The solver outputs the compliance which is combined with the volume fraction violation as a loss. The loss is then backpropagated to learn the weights of the topology network.
@@ 3.2 Multi-scale topology optimization with neural networks
The goal of using the neural network to represent the topology is to enforce compatibility between cells. The compatibility is achieved with the extension of the cell boundaries, which is illustrated in Figure 2. In 2D, we extend the boundary of each cell by 1.2 times to cover all its 8 neighboring cells. If rotation of the cell is required, the boundary is further extended to 1.6 times to cover all 8 neighboring cells. For the extended regions that are in neighboring cells, the domain coordinates use the local and global coordinates for neighboring cells. We use 4 node square elements for finite element analysis. Within the boundary, we sample the domain at uniform intervals that directly correspond with the finite element mesh. During training, the iterative updates to the design domain are carried out on these sampling points.
@@
As illustrated in Figure 3, the outer region for the 8 neighboring cells consists of the area highlighted in solid red lines and the outer region for the center cell is highlighted in blue dashed lines. The extent of the sampling for each cell is also expanded by 1.2 times. For the regular sampling for the continuous multi-scale structure, each local coordinate is normalized between -0.5 and 0.5. During the computation of the boundary loss, each cell is further expanded by 1.2 times such that the local coordinates are in the range of -0.6 to 0.6. During the optimization, the boundary loss is turned on after 50 epochs.
During optimization, the topology network outputs the density value at the center for each element. These density values are then sent to the finite element solver to calculate the homogenized stiffness tensor based on the SIMP interpolation. [I see this part could be controversial, especially mixing anisotropy with SIMP. You may want to acknowledge that there are other ways to do this, and explain why you did what you did.] We adopted the educational MATLAB code by Xia and Breitkopf [27] for inverse homogenization and programmed the FE code in Python to connect with the neural network that is implemented in TensorFlow.
The homogenization is based on an energy-based approach where unit test strains are directly imposed on the boundaries of the base cell. We note that , , and . A weight is assigned to each of the components of the stiffness tensor.
The finite element solver is treated as a black box within the neural network. It takes in the density of each element and outputs the homogenized stiffness tensor and the sensitivity for each element with respect to the stiffness tensor. Variables that are being optimized are the weights W and kernels K of the neural network. Adam[28] is used to train the neural network. The constrained optimization problem needs to be transferred into an unconstrained minimization problem for neural networks. The combined loss function is
The compliance is calculated for each cell and normalized against the compliance which is calculated from a homogenized density field with volume fraction of . is the penalty factor volume fraction of each cell . Starting from the first epoch, it gradually increases to 100 at the end of the optimization. is the penalty factor for the boundary loss which remains 0 until epoch 50 and gradually increases to 1 at the end of the optimization. We choose epoch 50 as we can observe some microstructure starts to appear while the domain is still largely not converged. One example of convergence history is plotted in Figure 5. A similar loss function for a single scale has also been used by Chandrasekhar and Suresh [6] for compliance minimization and volume fraction constraint with neural networks.
@@ 4.1 Multi-scale topology optimization with varying stiffness tensor
The example in Figure 4 consists of 8x8 cells on the coarse scale, 30x30 elements on the fine scale, and a uniform volume fraction target of 0.4. The stiffness tensor weights are configured such that in the horizontal direction is increased. In the vertical direction is increased. We run the optimization for 300 epochs and plot out a snapshot of the multi-scale structure during training in Figure 5. The optimization takes 9 minutes to run. During optimization, we observe that cells converge in clusters. On the top right corner of Figure 5 is a large cluster where the cells converged faster. As the optimization progresses, the front with converged microstructure gradually progresses to the bottom left. This could indicate that the neural network is using cells that are closer in global coordinates to interpolate.
@@
In addition to changing the stiffness tensor weights, the volume fraction can also be varied. Compared to the example in Figure 4, the example in Figure 6 has the volume fraction target increasing from 0.3 to 0.5 in the horizontal axis. We can observe the thickness of the multi-scale structure smoothly increased to accommodate for the changing volume fraction. The compliance evaluated in each cell is plotted in Figure 6 (a) where it is lower with a higher volume fraction.
@@
We then configure a problem with varying volume fraction and bulk modulus maximization. This allows us to compare the performance of the theoretical Hashin–Shrickman (HS) upper bound [29]. The converged multi-scale structure is plotted in Figure 7 (b). The volume fraction target in the center is 0.6 and is gradually decreased to 0.4 on the outer edges. We threshold the density field at 0.4. The bulk modulus of each cell and the corresponding HS upper bound are calculated with the thresholded density field. The ratio between the computed bulk modulus versus the HS upper bound is then plotted in 7 (a). On average, across 64 total cells and between 0.4 and 0.56 volume fraction, we achieved 84.9 of the HS upper bound for bulk modulus. This is slightly lower compared to the work by Sridhara et al.[10] where for single microstructure design by the neural network, they achieved 90 of the HS upper bound in the volume fraction range. Slightly higher restrictions on design freedom can be expected with the connectivity constraints posed. We note that no explicit symmetry constraint is applied to this example. The resulting structure demonstrated 2-fold rotation symmetry. While mirror symmetry is missing. This is attributed to the architecture of the neural network. Inputting global coordinates that are from the I and III or II and IV quadrants leads to the same result. For results where mirror symmetry is desired, explicit mirroring of the global coordinates needs to be performed. Previous works with neural network-based topology optimization also required such mirroring [6].
4.2 Multi-scale topology optimization with cell rotation
The example in 8 (a) consists of a rotated stiffness tensor target of . Given a target stiffness tensor field with rotation for example Figure 8 (a), we can either rotate the homogenized stiffness tensor or rotate the geometry. Since the input to the neural network consists of local and global coordinates, we can apply linear transformation matrices to the local coordinates. To rotate the geometry, a 2D rotation matrix can be multiplied on the local coordinates for each cell:
To rotate the stiffness tensor in Voigt notation requires the rotation matrix for rotating the basis of the stiffness tensor.
Where the rotation can be applied to the stiffness tensor with the following equation:
We can then compare the effect of rotating the geometry versus rotating the stiffness tensor. In Figure 8 (b), the geometry is rotated. In Figure 8 (c), the stiffness tensor is rotated. Rotating the geometry leads to microstructure better following the prescribed rotation. The transition is more abrupt in the case of rotating the stiffness tensor. The compliance reached under the same total number of epochs is lower with the rotating of the stiffness tensor.
We can also design multi-scale structures with negative Poisson’s ratio material with rotation. The Poisson’s ratio is defined as . We adopt the relaxed form of compliance objective function proposed in:
Where is the design epoch number and is chosen such that at the beginning of the optimization, the optimizer tends to maximize the horizontal and vertical stiffness. As the optimization progresses, the optimizer tends to minimize , such that materials with negative Poisson’s ratios are generated.
The result of an 88 coarse scale, 3030 element fine scale optimization with rotating microstructure is shown in Figure 9. We can observe that the multi-scale structure is well connected. And evaluating the Poisson’s ratio at each cell, as shown in Figure 9 (a), most of the cells, apart from the center 4 cells, have negative Poisson’s ratio.
4.3 Multi-scale topology optimization with distribution field
In this section, we delve into the utilization of stiffness tensor fields derived from coarse-scale topology optimization. While the previous subsection emphasized the application of smooth rotation and constant stiffness tensor fields, our focus now shifts toward more complicated stiffness tensor fields from coarse-scale topology optimization. The coarse scale topology optimization example shown in Figure 10 (a) consists of a single point load applied on the bottom right. The goal of the coarse-scale optimization is to obtain a principal strain field that can be used as a surrogate stiffness tensor field for subsequent fine-scale optimization. We run the coarse scale optimization at 2010 resolution with a volume fraction target of 0.6 under the assumption of isotropic material. We threshold the coarse scale design such that only pixels with a volume fraction higher than 0.2 are activated in the subsequent fine-scale optimization. The total number of active cells for this example is 132. We compute the principal strain and from the displacement and use the orientation to design the fine-scale structure. The principal strain is plotted in Figure 10 (b) where regions with similar orientation and maximum to minimum strain can be seen.
In the fine-scale optimization, we map the maximum and minimum principal strain to the stiffness tensor and as follows:
The resolution for each cell is 2020 with a uniform volume fraction target of 0.4 across all cells. We run a a total of 400 epochs which takes 17 minutes. The converged multi-scale structure is shown in Figure 10 (c). We can observe a good correlation between the fine-scale structure and the orientation of the stiffness field. The fine-scale structures show clustering behavior, while the transition between the clusters is smooth.

### v2

@@ Abstract
A long-standing challenge in multi-scale structural design is ensuring proper connectivity between the constituent cells as each cell is being optimized toward its theoretical performance limit. We propose a new method for multi-scale topology optimization that seamlessly ensures compatibility across neighboring microstructure cells without the need for a catalog of unit cells and associated connectivity rules. Our approach consists of a topology neural network that encodes each cell’s microstructure as well as the distribution of the optimized cells within the design domain as a continuous and differentiable density field. Each cell is optimized based on a prescribed elasticity tensor that also accommodates in-plane rotations. The neural network takes as input the local coordinates within a cell to represent the density distribution within a cell, as well as the global coordinates of each cell to generate spatially varying microstructure cells. As such, our approach models an n-dimensional multi-scale optimization problem as a 2n-dimensional inverse homogenization problem using neural networks. During the inverse homogenization of each unit cell, we extend the boundary of each cell by scaling the input coordinates such that the boundaries of neighboring cells are blended and thus jointly optimized. This enables inverse homogenization of the cells to be consistently connected without introducing discontinuities at the cell boundaries. We demonstrate our method through the design and optimization of several graded multi-scale structures.
@@ 1 INTRODUCTION
There has been a recent increase in machine learning-driven topology and multi-scale structures optimization approaches, particularly using neural networks for performing structural optimization. Both data-driven [1, 2, 3, 4, 5] and direct training-based [6, 7, 8, 9, 10, 11] approaches have been explored. Data-driven approaches typically require large training sets and long training times. They perform instant optimal topology generation during inference time. Direct training approaches use the neural network to represent the density field of a single or a group of designs. Prior works on direct training approaches focus on using the neural network with spatial coordinates as input and density as the output to represent the geometry [9, 10, 6]. Inverse homogenization refers to the optimization of microstructure with the desired property [12, 13]. Sridhara et al. [10] applied direct neural network-based inverse homogenization in designing single microstructure cells. Unfortunately, no direct neural network-based optimization for multi-scale structures has been developed. One key challenge from microstructure design to multi-scale structure is ensuring good connectivity between cells [13]. Connecting individual optimized microstructure cells into a multi-scale structure is called de-homogenization [13]. Cell rotation is included in de-homogenization in which cells can perform very close to the theoretical limit[13]. De-homogenization is a post-processing step to obtain a well-connected multi-scale design introduced by Pantz and Trabelsi [14]. During inverse homogenization, connectivity can also be encouraged where Garner et al. [15] combined adjacent cells and Liu et al. [16] formulated a constraint on the boundary of each cell. However, these previous methods did not incorporate rotation during inverse homogenization. On the other hand, de-homogenization relies on post-processing to add rotation. 111Example code is available at https://github.com/HongRayChen/MSNN
@@
In this work, we aim to use a neural network to represent the design of a multi-scale structure such that connectivity and rotation can be enforced. We configure a neural network to take in both local and global spatial coordinates to represent multi-scale structural designs. We focus on 2D multi-scale structures; hence, the coordinates are two-dimensional. The local coordinates represent the micro-scale microstructure. The global coordinates correspond with the macro-scale cell location. Given the local and global coordinate pairs as input, the neural network outputs the density for that specific location as a value between 0 and 1. During training222Here “training” refers to optimization and back-propagation. Training has been used in previous related works [6, 11] but this is different from data-driven machine learning, the microstructure for each cell is optimized to maximize or minimize a specific material property, derived from the elasticity tensor. We extend the unit cell boundary to its neighboring cells during inverse homogenization to encourage connectivity between cells. Since the inputs to the neural network are spatial coordinates, transformation matrices, for example, scaling, and rotation, can be applied to the coordinates. We apply rotation matrices to the local input coordinates such that the cell is oriented similarly to de-homogenization[14]. The difference is that in our approach the orientation of the cell is co-optimized during the training of the neural network instead of a post-processing step.
@@
accommodates rotation of cells with neural network geometry representation
@@ 2 RELATED WORK
Multi-scale topology optimization: Multi-scale structures offer the potential to deliver outstanding performance, being inherently lightweight, durable, and versatile. Bendsøe and Kikuchi [17] introduced the homogenization approach for topology optimization. The SIMP method [18, 19] considers the relative material density in each element of the Finite Element (FE) mesh as design variables, allowing for a simpler interpretation and optimized designs with more clearly defined features. A comprehensive review by Wu et al. [13] discusses the existing approaches of multi-scale topology optimization and analysis of their strength and applications. Research that targets improving the compatibility of multi-scale structures includes Garner et al. [15], where adjacent cells were combined when inverse homogenization is performed such that compatibility can be improved. However, adding neighboring cells to the optimization will increase the computation time for FE analysis. We instead extend the cell boundary into neighboring cells where the FE resolution remains the same. Liu et al. [16] formulated a constraint on the boundary of each cell such that given n types of desired microstructure, all these n types share the same boundary. This allows the optimized cell designs to be mutually compatible. One possible limitation is that the total types of cell designs are limited and as the number of cell designs increases, the restriction on the boundary can be increased. In this work, we establish connectivity between neighboring cells, where the limitation on geometry is less. Connectivity is also important when designing multi-scale structures given a material exemplar. Senhora et al. [20] use spinodal architected materials to place material along principal stress directions. A library of microstructure can be pre-generated and connectivity between cells can be improved through mapping [21] or inherent connectivity of the microstructure dataset [22]. However, the microstructure performance gamut can be limited by the generated dataset.
@@ 3.1 Neural network-based topology representation
The topology network (Figure 1), learns a density field differently as compared to typical topology optimization which represents the density field as a finite element mesh. The topology neural network takes in local domain coordinates , as well as the global coordinate for each cell. The global coordinates get concatenated with the local coordinates to form the input to the topology network, . The local coordinates are normalized between to . The global coordinates are normalized between to along the longest axis. Given the local and global coordinates, the topology network outputs the density value at each coordinate point. The domain coordinates represent the center of each element in the design domain. During topology optimization, a batch of domain coordinates that correspond to the mesh grid and the corresponding strain energy field is fed into the topology network. The output is then sent to the FE solver. The solver outputs the stiffness objective which is combined with the volume fraction violation as a loss. The loss is then backpropagated to learn the weights of the topology network.
@@
For a typical multi-scale problem, we employ around 5k frequency kernels and a same number of corresponding trainable weights. This means a total of 10k trainable parameters to represent a multi-scale structure.
3.2 Inverse Homogenization
We rely on an energy-based homogenization approach to compute the elasticity tensor of each cell and design the material distribution of each cell. A detailed review and accompanying code can be found in the work from Xia and Breitkopf [28]. Nevertheless, for the sake of clarity, we include a brief review of the formulation of homogenization and inverse homogenization in this section.
Homogenization aims to evaluate the equivalent constitutive behavior of repetitive microstructures. A compact and easy-to-implement formulation for homogenization is the energy-based approach based on the average stress and strain theorems. The homogenized stiffness tensor can be approximated in terms of the discretized N mutual energy of elements of the design domain .
Once the constitutive stiffness tensor is computed, inverse homogenization refers to the minimization of a combination of the homogenized stiffness tensor.
s.t. | ||||
The optimization equation involves , the global stiffness matrix, while and represent the global displacement vector and the external force vector for the test case , respectively. Here, refers to the spatial dimensionality, is the volume of an individual element, and defines the upper limit for the volume fraction. The objective function depends on the homogenized stiffness tensors. For example, the maximization of the bulk modulus can be written with an objective function:
@@ 3.3 Multi-scale topology optimization with neural networks
The goal of using the neural network to represent the topology is to enforce cell compatibility. The first mechanism to encourage compatibility is extending the cell boundaries during inverse homogenization, illustrated in Figure 2. In 2D, we extend the boundary of each cell by 1.2 times to cover all its 8 neighboring cells. If rotation of the cell is required, the boundary is further extended to 1.6 times to cover all 8 neighboring cells. For the extended regions that are in neighboring cells, the domain coordinates use the local and global coordinates for neighboring cells. We use 4 node square elements for finite element analysis. Within the boundary, we sample the domain at uniform intervals that directly correspond with the finite element mesh. During training, iterative updates to the design domain are carried out on these sampling points.
@@
As illustrated in Figure 3, the outer region for the 8 neighboring cells consists of the area highlighted in solid red lines and the outer region for the center cell is highlighted in blue dashed lines. The extent of the sampling for each cell is also expanded by 1.2 times. For the regular sampling for the continuous multi-scale structure, each local coordinate is normalized between -0.5 and 0.5. During the computation of the boundary loss, each cell is further expanded by 1.2 times such that the local coordinates are in the range of -0.6 to 0.6. We choose the extension ratio based on heuristics such that a small portion of the neighboring cells is covered while not being too big that most of the FE simulation patch consists of neighboring cells instead of the center cell. During the optimization, the boundary loss is turned on after 50 epochs.
During optimization, the topology network outputs the density value at the center for each element. These density values are then sent to the finite element solver to calculate the homogenized stiffness tensor based on the SIMP interpolation. We adopted the educational MATLAB code by Xia and Breitkopf [28] for inverse homogenization and programmed the FE code in Python to connect with the neural network that is implemented in TensorFlow.
The homogenization is based on an energy-based approach where unit test strains are directly imposed on the boundaries of the base cell. In order to achieve a different ratio of the stiffness tensor target, a weight is assigned to each of the components of the stiffness tensor.
The finite element solver is treated as a black box within the neural network. It takes in the density of each element and outputs the homogenized stiffness tensor and the sensitivity for each element with respect to the stiffness tensor. Variables that are being optimized are the weights W and kernels K of the neural network. Adam[29] is used to train the neural network with a learning rate of 0.002. The constrained optimization problem needs to be transferred into an unconstrained minimization problem for neural networks. The combined loss function is
The objective of stiffness tensor is calculated for each cell and normalized against the stiffness objective which is calculated from a homogenized density field with volume fraction of . is the penalty factor volume fraction of each cell . Starting from the first epoch, it gradually increases to 100 at the end of the optimization. is the penalty factor for the boundary loss which remains 0 until epoch 50 and gradually increases to 1 at the end of the optimization. We choose epoch 50 as we can observe some microstructure starts to appear while the domain is still largely not converged. One example of convergence history is plotted in Figure 5. A similar loss function for a single scale has also been used by Chandrasekhar and Suresh [6] for compliance minimization and volume fraction constraint with neural networks.
@@ 4.1 Multi-scale topology optimization with varying stiffness tensor
The example in Figure 4 consists of 88 cells on the macro-scale, 3030 elements on the micro-scale, and a uniform volume fraction target of 0.4. The stiffness tensor weights are configured such that in the horizontal direction is increased. In the vertical direction is increased. We run the optimization for 300 epochs and plot out a snapshot of the multi-scale structure during training in Figure 5. The optimization takes 9 minutes to run. During optimization, we observe that cells converge in clusters. On the top right corner of Figure 5 is a large cluster where the cells converged faster. As the optimization progresses, the front with converged microstructure gradually progresses to the bottom left. This could indicate that the neural network is using cells that are closer in global coordinates to interpolate.
@@
In addition to changing the stiffness tensor weights, the volume fraction can also be varied. Compared to the example in Figure 4, the example in Figure 6 has the volume fraction target increasing from 0.3 to 0.5 in the horizontal axis. We can observe the thickness of the multi-scale structure smoothly increased to accommodate for the changing volume fraction. The compliance evaluated in each cell is plotted in Figure 6 (a), where it is lower with a higher volume fraction.
@@
We then configure a problem with varying volume fraction and bulk modulus maximization. This allows us to compare the performance of the theoretical Hashin–Shrickman (HS) upper bound [30]. To maximize for the bulk modulus, the weightage on the stiffness tensor is configured as in equation 4. The converged multi-scale structure is plotted in Figure 7 (b). The volume fraction target in the center is 0.6 and is gradually decreased to 0.4 on the outer edges. We threshold the density field at 0.4. The bulk modulus of each cell and the corresponding HS upper bound are calculated with the thresholded density field. The ratio between the computed bulk modulus versus the HS upper bound is then plotted in 7 (a). On average, across 64 total cells and between 0.4 and 0.56 volume fraction, we achieved 84.9 of the HS upper bound for bulk modulus. This is slightly lower compared to the work by Sridhara et al.[10] where for single microstructure design by the neural network, they achieved 90 of the HS upper bound in the volume fraction range. Slightly higher restrictions on design freedom can be expected with the connectivity constraints posed. We note that no explicit symmetry constraint is applied to this example. The resulting structure demonstrated 2-fold rotation symmetry. While mirror symmetry is missing, this is attributed to the architecture of the neural network. Inputting global coordinates from the I and III or II and IV quadrants leads to the same result. For results where mirror symmetry is desired, explicit mirroring of the global coordinates needs to be performed. Previous works with neural network-based topology optimization also required such mirroring [6].
Two potential issues with the previous experiments are the linearly increasing memory consumption and longer runtime with full batches. Therefore, we explored mini-batches to reduce the number of cells seen at every backpropagation step and mini-epochs for which we only let the optimizer see a subset of cells every epoch to exploit the interpretability of neural networks. We compared runtime and the percentage of the Hashin–Shrickman upper bound achieved under different batch sizes and total epochs. In Figure 8 experiments (c, d), we used mini-batches where each batch contained 1/2 or 1/4 of the total cells. This setup resulted in stable convergence but slightly increased runtime due to more frequent backpropagation. In experiments (e, f), we tested whether the neural network’s interpretability allowed using a subset of cells per epoch, referred to as a mini-epoch. Since neighboring cells share similar global coordinates, we observed that during convergence, the neural network seems to leverage the knowledge of neighboring cells to infer the overall microstructure, effectively accelerating convergence. This behavior suggests that the network learns to generalize the spatial relationships and extrapolate the microstructure patterns from a subset of the data. Additionally, as neural networks are designed to approximate continuous functions, the learned representation appears to smooth out local variations and capture global trends, further enhancing training efficiency with reduced data. With 1/2 or 1/4 of the cells per mini-epoch, the runtime was reduced to nearly half or a quarter of the full-batch runtime, with only a slight trade-off in mechanical performance.
4.2 Multi-scale topology optimization
In this section, we delve into concurrent optimization of the two-scale structure. While the previous subsection emphasized the application of predefined stiffness tensors, our focus now shifts toward simultaneous optimization of the density distribution of the macro-scale structure and the micro-scale individual microstructure where similar optimization schemes are explored in [31, 32]. The neural network architecture remains the same as in the previous examples. We modify the loss function to include the compliance instead of the stiffness tensor target.
The multiscale topology optimization problem considered here assigns a distinct microstructure to each macro finite element. Let the macrostructure be discretized into elements, and for each macro element, we define a corresponding material microstructure discretized into elements. The macro design variables are denoted by , and the micro design variables (for each macro cell) by . The optimization problem is stated as:
Subject to: | ||||
Here, is the structural compliance. The macro topology is determined by , and each macro element’s effective properties depend on its own micro topology described by . imposes a limit on the macro material usage, and enforces a limit on the micro material consumption. The lower bounds and avoid numerical singularities. The displacement field at the macro scale is denoted by , with virtual displacement . The operator is a bilinear form representing the internal energy, and is a linear form associated with external forces. At the macro scale, the principle of virtual work leads to:
In the above, is the body force, and is the boundary traction on the Neumann boundary . The tensors and are derived via a material interpolation scheme akin to the modified SIMP approach, given by:
where is the base material elasticity tensor, prevents numerical singularity, and is the penalization exponent. The homogenized tensor depends on the local micro design:
A variety of gradient-based methods may be employed to solve this formulation. Thus, the sensitivities (first-order derivatives) of and the volume constraints , with respect to the design variables are needed. The partial derivatives of the objective and macro volume constraint with respect to read:
Likewise, the first-order derivatives of and with respect to the micro design are expressed as:
The derivative of the homogenized elastic tensor with respect to is given by:
With these sensitivities, standard gradient-based optimization procedures can be used to simultaneously update the macro design variables and each element’s corresponding micro design variables . In implementation, the changes occur in implementing a multiscale FE solver, which computes the compliance for the entire structure and adds the volume fraction constraint for the macro scale structure. The combined loss function for concurrent multiscale topology optimization is:
Furthermore, the macro-scale design is represented by a second neural network with the domain coordinate input reduced to 2D: . While the frequency kernels K and W are reduced to align with the 2D input.
We demonstrate the multiscale topology optimization with three examples. The boundary condition is shown in Figure 11 (a-c). The converged multi-scale structure is shown in Figure 11 (d-f). One benefit of using a similar neural network to represent the macro scale design field is that it allows us to upsample as well. We use the upsampled macro scale density field as a filter to threshold the microstructures. Compared to filtering with the grid-based macro-scale design field, the continuous macro-scale density field allows smooth boundaries to be extracted that better follow microstructures. We observe that good cell compatibility is achieved in the multi-scale structure.
Similar to the ”mini-epoch” example in the previous subsection, we also demonstrate the neural network’s interpolation ability with multi-scale topology optimization. We configure a 1/4 mini-epoch, which uses four microstructure cells per macro element. In Figure 12 (a), we choose one out of four microstructure cells every epoch to conduct homogenization and optimization. We cycle through each cell every four epochs. We run the same number of total epochs and the same boundary condition as the previous example. We observe that the converged result also consists of good cell connectivity with no significant increase in runtime.
4.3 Metamaterial system design
Metamaterials are engineered materials whose unique properties stem from the arrangement of their microstructures rather than the materials they are made of [33]. Mechanical metamaterials are a specific type designed to achieve mechanical characteristics that cannot be attained through their base material alone. Prior work by Wang et al. demonstrated metamaterial design of metamaterials to match a target displacement through the microstructure selection with a deep generative model [34]. We are motivated by prior work in designing metamaterials to match a target displacement with our neural network. For the metamaterial design problem, similar to the prior work by Wang et al., we limit the scope to microstructure design. One difference is that since we are designing the microstructure real-time, we also need an additional optimization objective of bulk modulus maximization to ensure the result is overall structurally sound. Otherwise, with only the displacement target as the objective, the result may lack structural performance in other loading conditions.
The objective function first consists of a target displacement where the displacement u of the multiscale structure is trying to match. Same with Wang et al., we minimize the difference in displacement with an L2 loss:
Following Wang et al., we only design the microstructure in metamaterial system design. Thus we modify their sensitivity function to compute the derivative of the objective function w.r.t micro design field :
The combined loss function also resembles the previous examples of bulk modulus minimization, but it includes an additional term for displacement mismatch.
Specifically for displacement matching, is a vector with the value one at the degrees of freedom corresponding to the displacement targets and with zeros everywhere else.
We configure two examples with different boundary conditions Figure 13 (a)(b). The target displacement is highlighted in red in each boundary condition figure. In the first example, the boundary condition and displacement target are similar to one of the examples in Wang et al. [34]. A load is applied on the right-hand side with a target displacement along the center of the structure. We configure the second example with load applied to the bottom center and target displacement prescribed to the top surface with a region of enforced solid highlighted in gray. Some potential applications for the second boundary condition include flexible haptics devices and jet engine intake ramps. We plot out the target and achieved displacement in Figure 13 (c)(d), where we observe minimal errors. The converge metamaterial is shown in Figure 13 (e)(f) with deformed metamaterial shown in Figure 13 (g)(h).